An asset owned by Photon Environmental was book depreciated by the Straight Line method over a 5-year period with book values of $260419.8 and $202412.2 in years 2 and 3, respectively. Calculate the unadjusted base value (B) based on two given years book value using the Straight Line method of depreciation.
Answer #1

Useful life=n=5 years

Deprecation rate=100%/5=20%

Book value after 2nd year=BV2=$260419.80

Book Value after 3rd year=BV3=$202412.20

Depreciation=BV2-BV3=260419.80-202412.20=$58007.60

We know that depreciation value is same for all years in case of Straight Line depreciation.

Book Value at the end of year 5=BV3-2*D=202412.20-2*58007.60=$86397

We know that Book value after useful life is equal to salvage, S.

Let the first cost i.e. unadjusted base value be B

We know D=(B-S)*20%

58007.60=(B-86397)*20%

290038=B-86397

B=290038+86397=$376435